OpenStudy (anonymous):

Please help! You want to do an experiment using 47Ca (half-life=4.53 days). You will need 6.00 μg of the nuclide. What mass of CaCO3 (100% enriched in 47Ca) must you order if it will take 49 hours to be delivered to you from the nuclear facility where it is prepared?

OpenStudy (aaronq):

use this equation: \(ln[A]_t=-\dfrac{ln(2)*t}{t_{1/2}}+[A]_o\) \([A]_o\)=initial amount of substance t=time \(t{1/2}\)=half-life \([A]_t\)= is the amount of substance after t time has passed

OpenStudy (anonymous):

Ah...How do you calculate for t 1/2?

OpenStudy (aaronq):

It's given in the question, half-life=4.53 days

OpenStudy (anonymous):

Oh okay! Thank you.

OpenStudy (aaronq):

no problem, good luck!

OpenStudy (anonymous):

Also, will the answer be for the Ca or for CaCO3?

OpenStudy (aaronq):

Ca only

OpenStudy (anonymous):

Ah...Do I use molar mass and convert the amount into CaCO3?

OpenStudy (anonymous):

One more question (Sorry DX): What will be the units then?

OpenStudy (aaronq):

we'll what you'll find with the equation above will be the mass of Ca you need. Convert that to moles, which is the same as CaCO3, multiply those moles by the molar mass of CaCO3. This is in the units that you started with, (from half-life) days, (from mass) \(\mu g\)

OpenStudy (anonymous):

Okay. I'll check in once I figure out the answer and see if it's right xD

OpenStudy (aaronq):

make sure you use the same units for t and \(t_{1/2}\), either both in hours or both in days

OpenStudy (anonymous):

Yep, did so.

OpenStudy (aaronq):

good stuff

OpenStudy (anonymous):

Okay...this is a lot of steps, and now, I'm really confused. X _ X

OpenStudy (aaronq):

lol focus on finding the mass of Ca you need initially.

OpenStudy (anonymous):

Found it ; ~ ;

OpenStudy (aaronq):

sweet. now, convert it to moles of Ca. The mass needs to be in grams

OpenStudy (anonymous):

Got it.

OpenStudy (aaronq):

now since \(n_{Ca}=n_{CaCO_3}\), multiply \(n_{Ca}\) by the molar mass of CaCO3. you're basically using this formula \(n=\dfrac{m}{M}\)

OpenStudy (anonymous):

n?

OpenStudy (anonymous):

???

OpenStudy (aaronq):

n is moles

OpenStudy (anonymous):

Got it.....

OpenStudy (aaronq):

good job

OpenStudy (anonymous):

That's it?

OpenStudy (aaronq):

The question asked for the mass of CaCO3, right? that's the last thing you found

OpenStudy (anonymous):

Didn't get it xP

OpenStudy (anonymous):

Is there an easier way to do it?

OpenStudy (aaronq):

hm no, double-check your units or post what you did i'll tell you where you went wrong.

OpenStudy (anonymous):

I got 9.29 picograms of calcium.

OpenStudy (aaronq):

\(\mu g =micrograms=1.0*10^{-6}g\)

OpenStudy (anonymous):

Oops. x_x

OpenStudy (aaronq):

maybe thats where it went sour

OpenStudy (anonymous):

2.10416 micrograms

OpenStudy (anonymous):

and I got 5.25016 x 10^-8 mol :3

OpenStudy (aaronq):

hm ill check it, give me a min

OpenStudy (anonymous):

Sure. Thanks~

OpenStudy (aaronq):

yep, it's right. now multiply that by the molar mass of CaCO3

OpenStudy (anonymous):

The amount of mol of Ca by the molar mass? Or do I have to keep it in g?

OpenStudy (aaronq):

nope. the molar mass is in units of g/mol by multiplying moles * molar mass, you eliminate the "/mol" units, and end up with units of "grams" only.

OpenStudy (anonymous):

Gotcha.

OpenStudy (anonymous):

Alright. I got 5.2547 x 10^-6 g

OpenStudy (aaronq):

yep, thats it

OpenStudy (anonymous):

Let me try the answer in my program...

OpenStudy (aaronq):

they might want you to implement significant digits and round

OpenStudy (anonymous):

They want me to put down a decimal, and apparently, it's still wrong x _ x

OpenStudy (aaronq):

then it would be 5.3 x 10^-6 g

OpenStudy (anonymous):

Still would accept more digits x - x It'd let the answer through and say it's correct :3

OpenStudy (aaronq):

hm k, it should be right if it wants 1 decimal place

OpenStudy (anonymous):

I'm trying a different way...Let me see if it's right x - x